OpenStudy (anonymous):

An object is seen at a position (0, -10, -2)m after moving for 12s with an avg velocity of (-1/4, -1, -1/2) m/s. What was the magnitude of the displacement of the object during those 12s (I got (-3,-12,-6)? What was the initial location before those 12s (I got 3,2,4). If the object speeds up by 40% where will it end up in 12 more seconds (Would I multiply the given position by the product of 12 and 1.4?)

OpenStudy (anonymous):

Am I correct? I'm confused now because I think I may have done something wrong!

OpenStudy (anonymous):

Do you have to use fractions? -.25, -1, -.5 m/s *12 = -3, -12, -6. (0,-10,-2) -(-3, -12 -6)= 3,2,4, multipy the velocity times 1.4, figure new magnitude, and new end position. -.35, -1.4, -.7 *12= ....

OpenStudy (anonymous):

Oh okay! That's what I was thinking... Thanks @Fixitupman!